# On silting complexes associated to $n$-silting modules

### v1

@@ 3.3. More general construction
Here, note that since $A_{\mathfrak{p}}$ is Cohen–Macaulay for any non-maximal prime $\mathfrak{p}$, we have $\operatorname{ht}(\mathfrak{p})=\operatorname{depth}(A_{\mathfrak{p}})$ and so the cosilting complex in $\mathbf{D}(A_{\mathfrak{p}})$ corresponding to $f_{\mathfrak{p}}$ via Theorem 3.1 indeed represents a $\operatorname{ht}(\mathfrak{p})$-cotilting module $C(\mathfrak{p})$. Taking the standard injective resolution of $D_{\widehat{A}}$, we know that $\Omega$ is quasi-isomorphic to a complex of injectives concentrated in degrees between 0 and $d$. Furthermore, $\Omega$ is a cogenerator of $\mathbf{D}(A)$, as it is cosupported on the whole spectrum, recall here from [18, p. 282, 287] that $\mathbf{R}{\rm Hom}_{A}(k,D_{\widehat{A}})\cong k[-d]$.
Using Lemma 1.1, it thus suffices to show that $\Omega^{\varkappa}\in{}^{\perp_{>0}}\Omega$ for any set $\varkappa$. Recall from [23, §5] that ${\rm Hom}_{\mathbf{D}(A)}(Y,D_{\widehat{A}}[i])=0$ if and only if $H_{\mathfrak{m}}^{d-i}(Y)=0$. Thus,
@@
It thus follows easily that $^{\perp_{>0}}\Omega=\mathcal{V}_{f}$. Clearly, $C(\mathfrak{p})\in\mathcal{V}_{\mathfrak{p}}\subseteq\mathcal{V}$ for any prime ideal $\mathfrak{p}$ other than $\mathfrak{m}$. It remains to check that $D_{\widehat{A}}\in\mathcal{V}$, or that $\operatorname{depth}D_{\widehat{A}}\geq d$, but that follows by the aforementioned isomorphism $\mathbf{R}{\rm Hom}_{A}(k,D_{\widehat{A}})\cong k[-d]$. This shows that $\Omega$ is an $(n+1)$-term cosilting complex. By the assumptions on $A$, we know that $H_{\mathfrak{m}}^{i}(A)=0$ for all $0<i<d$, but then also $H^{i}(D_{\widehat{A}})=0$ for all $0<i<d$, as $D_{\widehat{A}}\cong{\rm Hom}_{A}(\mathbf{R}\Gamma_{V(\mathfrak{m})}(A)[d],E(\mathfrak{m}))$, see e.g. [23, (5.8)]. It follows that $\Omega$ represents a $d$-cosilting module.
A monotone perversity $f:\operatorname{Spec}(A)\to\mathbb{Z}$ is called comonotone if $f(\mathfrak{q})\leq f(\mathfrak{p})+1$ whenever $\mathfrak{p}\subsetneq\mathfrak{q}$ are prime ideals such that there is no prime ideal $\mathfrak{o}$ such that $\mathfrak{p}\subsetneq\mathfrak{o}\subsetneq\mathfrak{q}$. These special perversities, under mild conditions (such as the existence of a dualizing complex) on $A$, correspond precisely to t-structures in the bounded derived category $\mathbf{D}^{b}(\mathrm{mod}(A))$ of finitely generated $A$-modules, see [5], [36], [35]. By [31, Corollary 6.17], the corresponding cosilting t-structures are cotilting in this situation.
@@ 3.4. Failure of finite type for $n$-silting modules
Let now $T$ be the $2$-silting module corresponding to the $3$-term tilting complex $\Sigma$, which in turn corresponds to the function $f$ assigning $2$ to $\mathfrak{m}$ and 0 to all other prime ideals, as per Example 3.8. Let $\mathfrak{p}$ be a non-maximal prime ideal of $A$. Then $\Sigma_{\mathfrak{p}}$ is isomorphic to a projective $A_{\mathfrak{p}}$-module in degree 0 for any non-maximal prime $\mathfrak{p}$, see [21, Lemma 6.3]. It follows by adjunction that for any $A_{\mathfrak{p}}$-module $N$ we have $N[i]\in\Sigma^{\perp_{>0}}$ for all $i>0$. In particular, we have $W\in\Sigma^{\perp_{>0}}$, where $W=\prod_{\mathfrak{p}\neq\mathfrak{m}}E(A/\mathfrak{p})$ is the product of isoclasses of indecomposable injective $A$-modules not supported on the closed point $\mathfrak{m}$.
@@
Thus, the complex $\sigma$ witnesses tha $k^{n}$ is a 3rd syzygy of the zero cohomology module $M:=H^{0}(\sigma)$. We claim that this implies that the residue field $k$ itself is a 3rd syzygy (of some finitely generated $A$-module). Note that this is not a trivial fact, as $n$th syzygies are in general not closed under direct summands. In fact, [6, Theorem A] shows that over $A$, every non-trivial 33rd syzygy has $k$ as a direct summand. Once the claim is established, the desired contratiction follows directly from [28, Corollary 2.13], as $A$ is manifestly not a Gorenstein ring.
Set $B=A/k$, where $k$ denotes the 1-dimensional socle of $A$ generated by $x$. Following the argument of [28, Lemma 2.8], we have an exact sequence of the form
@@
The map $\pi$ is the induced projection and its kernel is easily checked using the Snake Lemma to be isomorphic to $B^{n-1}$. It follows that $L$ is an extension of $B^{n-1}$ by $N$, both of which are syzygy modules. By [16, Theorem B], (1st) syzygy modules are closed under extensions for the ring $A$, so that $L$ itself is a syzygy module, which in turn makes $k$ a 3rd syzygy. Indeed, the assumptions of loc. cit. are satisfied, as $A$ is Cohen–Macaulay at any non-maximal prime, $A$ admits a unique minimal prime ideal $(x)$, and the localization $A_{(x)}\cong k[[x,y^{\pm 1},z^{\pm 1}]]/(x^{2})$ is an artinian hypersurface, and thus is Gorenstein.

### v2

@@ 3.3. More general construction
Here, note that since $A_{\mathfrak{p}}$ is Cohen–Macaulay for any non-maximal prime $\mathfrak{p}$, we have $\operatorname{ht}(\mathfrak{p})=\operatorname{depth}(A_{\mathfrak{p}})$ and so the cosilting complex in $\mathbf{D}(A_{\mathfrak{p}})$ corresponding to $f_{\mathfrak{p}}$ via Theorem 3.1 indeed represents a $\operatorname{ht}(\mathfrak{p})$-cotilting module $C(\mathfrak{p})$. Taking the standard injective resolution of $D_{\widehat{A}}$, we know that $\Omega$ is quasi-isomorphic to a complex of injectives concentrated in degrees between 0 and $d$. Furthermore, $\Omega$ is a cogenerator of $\mathbf{D}(A)$, as it is cosupported on the whole spectrum, recall here from [19, p. 282, 287] that $\mathbf{R}{\rm Hom}_{A}(k,D_{\widehat{A}})\cong k[-d]$.
Using Lemma 1.1, it thus suffices to show that $\Omega^{\varkappa}\in{}^{\perp_{>0}}\Omega$ for any set $\varkappa$. Recall from [24, §5] that ${\rm Hom}_{\mathbf{D}(A)}(Y,D_{\widehat{A}}[i])=0$ if and only if $H_{\mathfrak{m}}^{d-i}(Y)=0$. Thus,
@@
It thus follows easily that $^{\perp_{>0}}\Omega=\mathcal{V}_{f}$. Clearly, $C(\mathfrak{p})\in\mathcal{V}_{\mathfrak{p}}\subseteq\mathcal{V}$ for any prime ideal $\mathfrak{p}$ other than $\mathfrak{m}$. It remains to check that $D_{\widehat{A}}\in\mathcal{V}$, or that $\operatorname{depth}D_{\widehat{A}}\geq d$, but that follows by the aforementioned isomorphism $\mathbf{R}{\rm Hom}_{A}(k,D_{\widehat{A}})\cong k[-d]$. This shows that $\Omega$ is an $(n+1)$-term cosilting complex. By the assumptions on $A$, we know that $H_{\mathfrak{m}}^{i}(A)=0$ for all $0<i<d$, but then also $H^{i}(D_{\widehat{A}})=0$ for all $0<i<d$, as $D_{\widehat{A}}\cong{\rm Hom}_{A}(\mathbf{R}\Gamma_{V(\mathfrak{m})}(A)[d],E(\mathfrak{m}))$, see e.g. [24, (5.8)]. It follows that $\Omega$ represents a $d$-cosilting module.
A monotone perversity $f:\operatorname{Spec}(A)\to\mathbb{Z}$ is called comonotone if $f(\mathfrak{q})\leq f(\mathfrak{p})+1$ whenever $\mathfrak{p}\subsetneq\mathfrak{q}$ are prime ideals such that there is no prime ideal $\mathfrak{o}$ such that $\mathfrak{p}\subsetneq\mathfrak{o}\subsetneq\mathfrak{q}$. These special perversities, under mild conditions (such as the existence of a dualizing complex) on $A$, correspond precisely to t-structures in the bounded derived category $\mathbf{D}^{b}(\mathrm{mod}(A))$ of finitely generated $A$-modules, see [6], [2], [36]. By [32, Corollary 6.17], the corresponding cosilting t-structures are cotilting in this situation.
@@ 3.4. Failure of finite type for $n$-silting modules
Let now $T$ be the $2$-silting module corresponding to the $3$-term tilting complex $\Sigma$, which in turn corresponds to the function $f$ assigning $2$ to $\mathfrak{m}$ and 0 to all other prime ideals, as per Example 3.8. Let $\mathfrak{p}$ be a non-maximal prime ideal of $A$. Then $\Sigma_{\mathfrak{p}}$ is isomorphic to a projective $A_{\mathfrak{p}}$-module in degree 0 for any non-maximal prime $\mathfrak{p}$, see [22, Lemma 6.3]. It follows by adjunction that for any $A_{\mathfrak{p}}$-module $N$ we have $N[i]\in\Sigma^{\perp_{>0}}$ for all $i>0$. In particular, we have $W\in\Sigma^{\perp_{>0}}$, where $W=\prod_{\mathfrak{p}\neq\mathfrak{m}}E(A/\mathfrak{p})$ is the product of isoclasses of indecomposable injective $A$-modules not supported on the closed point $\mathfrak{m}$.
@@
Thus, the complex $\sigma$ witnesses tha $k^{n}$ is a 3rd syzygy of the zero cohomology module $M:=H^{0}(\sigma)$. We claim that this implies that the residue field $k$ itself is a 3rd syzygy (of some finitely generated $A$-module). Note that this is not a trivial fact, as $n$th syzygies are in general not closed under direct summands. In fact, [7, Theorem A] shows that over $A$, every non-trivial 33rd syzygy has $k$ as a direct summand. Once the claim is established, the desired contratiction follows directly from [29, Corollary 2.13], as $A$ is manifestly not a Gorenstein ring.
Set $B=A/k$, where $k$ denotes the 1-dimensional socle of $A$ generated by $x$. Following the argument of [29, Lemma 2.8], we have an exact sequence of the form
@@
The map $\pi$ is the induced projection and its kernel is easily checked using the Snake Lemma to be isomorphic to $B^{n-1}$. It follows that $L$ is an extension of $B^{n-1}$ by $N$, both of which are syzygy modules. By [17, Theorem B], (1st) syzygy modules are closed under extensions for the ring $A$, so that $L$ itself is a syzygy module, which in turn makes $k$ a 3rd syzygy. Indeed, the assumptions of loc. cit. are satisfied, as $A$ is Cohen–Macaulay at any non-maximal prime, $A$ admits a unique minimal prime ideal $(x)$, and the localization $A_{(x)}\cong k[[x,y^{\pm 1},z^{\pm 1}]]/(x^{2})$ is an artinian hypersurface, and thus is Gorenstein.